A large conducting plane has a surface charge density 1.0 × 10^-4 C/m^2. Find the electrostatic energy stored in a cubical volume of edge 1.0 cm in front of the plane.

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. σ = 1 * 10^-4 c/m^2 A = 1 cm = 1 * 10^-2 m a^3 = 10^-6 m The energy stored in the plane = 1/2 σ^2/ε base 0 = 1/2 91* 10^-4)^2/8.85 * 10^-12 = 10^4/17.7 = 564.97 The necessary electro static energy stored in a cubical volume of edge 1 cm infront of the plane = 1/2 σ^2/ε base 0 a^3 = 265 * 10^-6 = 5.65 = 10^-4 J